# Model-agnostic Coreset Selection via LLM-based Concept Bottlenecks

### v1

@@ C.1 Visualizing easy/challenging samples based on concept-based score
Here we present the algorithm for sampling the training examples to form the coreset based on the coverage-based selection methodology proposed by (Zheng et al., 2022). A crucial component of the algorithm is the cutoff rate $\beta$ which controls how many challenging samples should be removed from consideration when selecting the coreset. This is done to eliminate misclassified samples from the dataset since they can hurt the performance of the model trained on coreset, especially at high pruning rates. Previous works (Zheng et al., 2022, 2024) ablate the values of this cutoff ratio by training the downstream model on a range of values. In our work, we simply use the values proposed by the previous works and find that they work well for our score as well. The cutoff rates $\beta$ for different pruning rates $\alpha$ are as follows ($\alpha$, $\beta$). For CIFAR-10: (30%, 0), (50%, 0), (70%, 10%), (90%, 30%), for CIFAR-100: (30%, 10%), (50%, 20%), (70%, 20%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%). We used CCS for label-free CS as well and the cutoff rates used were for CIFAR-10: (30%, 0), (50%, 0), (70%, 20%), (90%, 40%), for CIFAR-100: (30%, 0), (50%, 20%), (70%, 40%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%).
Algorithm 1 Coverage-centric Coreset Selection (CCS) (Zheng et al., 2022)
@@
Table 6: Performance of concept-based coreset selection for emotion recognition task on Affectnet dataset. Coresets selected by our approach results in superior performance (higher F1 score) compared to Random for standard CS
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
EfficientNet Random $0.607_{\tiny\pm 0.006}$ $0.573_{\tiny\pm 0.025}$ $0.517_{\tiny\pm 0.006}$ $0.347_{\tiny\pm 0.055}$
Ours $0.603_{\tiny\pm 0.006}$ $0.577_{\tiny\pm 0.021}$ $0.537_{\tiny\pm 0.015}$ $0.450_{\tiny\pm 0.035}$
Table 7: Performance of concept-based coreset selection for emotion recognition task on BloodMNIST dataset. Coresets selected by our approach results in superior accuracy compared to Random for standard CS; highlighting our method’s effectiveness across various tasks.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
ResNet-18 Random $94.99_{\tiny\pm 0.27}$ $94.79_{\tiny\pm 0.18}$ $91.71_{\tiny\pm 0.29}$ $86.50_{\tiny\pm 0.66}$
Ours $94.64_{\tiny\pm 0.60}$ $94.06_{\tiny\pm 0.10}$ $92.26_{\tiny\pm 0.19}$ $87.44_{\tiny\pm 1.46}$
C.3 Concept-based CS for emotion recognition and biomedical image recognition
To validate the generalizability of our concept-based coreset selection method beyond image recognition tasks, we apply our concept-based CS approach to the task of emotion recognition and biomedical image recognition.
For emotion recognition, we use the Affectnet dataset (Mollahosseini et al., 2017) for our experiments. AffectNet is a large-scale facial expression dataset designed for training and evaluating affective computing models (Wang et al., 2022). It contains facial images collected from the internet using web search queries for emotion-related keywords in multiple languages. Each image is manually annotated for eight discrete emotion categories: neutral, happiness, sadness, surprise, fear, disgust, anger, contempt. For our experiments, we utilize an openly available version of this dataset  [^1] https://www.kaggle.com/datasets/noamsegal/affectnet-training-data, containing roughly $16000$ training and $14000$ testing samples.
According to our approach we first use LLaVA to extract concepts for the 8 emotion classes, using the following prompt, “What are the facial features that distinguish emotion class name from other emotion types. Focus on changes in eyes, nose, lips, eyebrows, mouth. Give the output separated by commas in a line.”. We get $5-10$ distinctive facial feature concepts for every emotion, for instance for emotion class happy, we get the following concepts, “wide open eyes”, “sparking eyes”, “smiling lips”, “open mouth”, “raised eyebrows”, “flushed cheeks”, “teeth barred”.
To test coreset performance, we use the EfficientNet model (Tan & Le, 2019) and report F1 scores for our coresets in Table 6. When compared against randomly selected coresets for the various pruning ratios, coresets selected via our concept-based approach achieve better performance.
For biomedical image recognition, we use the BloodMNIST (Acevedo et al., 2020) dataset from the MedMNIST (Yang et al., 2021, 2023a) which comprises of images of normal blood cells, captured from individuals without infection, hematologic or oncologic disease and free of any pharmacologic treatment at the moment of blood collection. It consists of a total of $17,092$ images and is organized into 8 classes (basophil,eosinophil,erythroblast,immature granulocytes(myelocytes, metamyelocytes and promyelocytes),lymphocyte,monocyte,neutrophil,platelet).
C.4 Experimental details
@@
For generating the importance score we pre-compute the concept similarity scores for the entire dataset and then train the concept-bottleneck layer (in block 2 of Fig. 1) for $100$ epochs across all experiments. This training only requires 800 seconds for Imagenet which is significantly more efficient than training the ResNet-34 model on Imagenet (requires roughly 8 hours without FFCV and about 4 hours with FFCV on two A-100 GPUs).
@@
After the coresets are selected, we use the setting and code from (Zheng et al., 2022) for training a ResNet-18 model for $40000$ iterations with a batch size of $256$ on the coresets for all pruning ratios for CIFAR-10/CIFAR-100. For Imagenet, we train ResNet-18, ResNet-34, and ResNet-50 models for $100$ epochs on the coresets identified by our method using the training code based on FFCV (Leclerc et al., 2023).
@@
The performance of the label-free CS is dependent on the quality of the pseudo-labels. Compared to the clustering-based approach used by ELFS (Zheng et al., 2024), our approach of using the zero-shot classification ability of CLIP models yields significantly better pseudo-label quality along with being simpler and more efficient to compute. Specifically, for CIFAR-10/100, pseudo-labels of the training set are computed using the CLIP L-14 model trained on the DataComp-1B dataset (Ilharco et al., 2021) yields an accuracy of $98.52$% and $87.28$% whereas for Imagenet it achieves an accuracy of $79.47$% which are better than the best pseudo-label accuracy obtained by the clustering approach in ELFS ($92.5$% and $66.3$% on CIFAR-10/100 and $58.8$% on Imagenet).

### v2

@@ C.5 Algorithm for stratified sampling using CCS [86]
Here we present the algorithm for sampling the training examples to form the coreset based on the coverage-based selection methodology proposed by [86]. A crucial component of the algorithm is the cutoff rate $\beta$ which controls how many challenging samples should be removed from consideration when selecting the coreset. This is done to eliminate misclassified samples from the dataset since they can hurt the performance of the model trained on coreset, especially at high pruning rates. Previous works [86, 87] ablate the values of this cutoff ratio by training the downstream model on a range of values. In our work, we simply use the values proposed by the previous works and find that they work well for our score as well. The cutoff rates $\beta$ for different pruning rates $\alpha$ are as follows ($\alpha$, $\beta$). For CIFAR-10: (30%, 0), (50%, 0), (70%, 10%), (90%, 30%), for CIFAR-100: (30%, 10%), (50%, 20%), (70%, 20%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%). We used CCS for label-free CS as well and the cutoff rates used were for CIFAR-10: (30%, 0), (50%, 0), (70%, 20%), (90%, 40%), for CIFAR-100: (30%, 0), (50%, 20%), (70%, 40%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%).
Algorithm 1 Coverage-centric Coreset Selection (CCS) [86]
@@
C.6 Experimental details
@@
For generating the importance score we pre-compute the concept similarity scores for the entire dataset and then train the concept-bottleneck layer (in block 2 of Fig. 2) for $100$ epochs across all experiments. This training only requires 800 seconds for Imagenet which is significantly more efficient than training the ResNet-34 model on Imagenet (requires roughly 8 hours on two A-100 GPUs). The accuracies of the models trained on the entire training set are $95.44$% and $78.74$% for ResNet(RN)-18 on CIFAR-10/100 and $72.4$% for RN-18, $75$% for RN-34, and $78.4$% for RN-50 on Imagenet.
@@
After the coresets are selected, we use the setting and code from [86] for training a ResNet-18 model for $40000$ iterations with a batch size of $256$ on the coresets for all pruning ratios for CIFAR-10/CIFAR-100. For Imagenet, we train ResNet-18, ResNet-34, and ResNet-50 models for $100$ epochs on the coresets identified by our method using the training code based on FFCV [37].
@@
The performance of the label-free CS is dependent on the quality of the pseudo-labels. Compared to the clustering-based approach used by ELFS [87], our approach of using the zero-shot classification ability of CLIP models yields significantly better pseudo-label quality along with being simpler and more efficient to compute. Specifically, for CIFAR-10/100, pseudo-labels of the training set are computed using the CLIP L-14 model trained on the DataComp-1B dataset [27] yields an accuracy of $98.52$% and $87.28$% whereas for Imagenet it achieves an accuracy of $79.47$% which are better than the best pseudo-label accuracy obtained by the clustering approach in ELFS ($92.5$% and $66.3$% on CIFAR-10/100 and $58.8$% on Imagenet).
For training the concept bottleneck layer we minimized the cross entropy loss using SGD with a learning rate of 1E-3, momentum of $0.9$ and a weight decay of 5E-4 for $100$ epochs.